Gdub08:

If a x b = 2a - 3b + ab, then 3 x 5 + 5 x 3 is equal to?

Gdub08:

I'm stuck

SnowyBreaks:

3x5+5x3=30

Gdub08:

@snowybreaks wrote:
3x5+5x3=30
but what do I do with the variables?

SnowyBreaks:

One sec let me work it out

SnowyBreaks:

a * b = 2a - 3b + ab 3 * 5 = 2(3) - 3(5) + (3)(5) = 6 - 15 + 15 = 6 5 * 3 = 2(5) - 3(3) + (5)(3) = 10 - 9 + 15 = 1 + 15 = 16 3 * 5 + 5 * 3 = 6 + 16 = 22 should be it

Gdub08:

@snowybreaks wrote:
a * b = 2a - 3b + ab 3 * 5 = 2(3) - 3(5) + (3)(5) = 6 - 15 + 15 = 6 5 * 3 = 2(5) - 3(3) + (5)(3) = 10 - 9 + 15 = 1 + 15 = 16 3 * 5 + 5 * 3 = 6 + 16 = 22 should be it
OHHHHHH!!!. Do I need to use Pemdas?

SnowyBreaks:

yes

Gdub08:

@snowybreaks wrote:
yes
Ok. I get it now. Thank you so much.

SnowyBreaks:

No problemm

SnowyBreaks:

Use this for an example later in time